# thanks: 0000-0002-5209-872Xthanks: 0000-0003-2028-6782

### v1

@@ IV.1 Solar Neutrinos
Nighttime solar neutrinos, contain a regeneration effect. It is known, based on oscillation probability parameters and some preliminary data, that the solar $\nu_{e}$ disappearance probability is expected to be slightly higher, on average, at night than during the day. While the probability oscillates quite quickly, directly observing those oscillations requires energy resolution and statistics far beyond what is achievable. Nonetheless, the averaged effect modifies the probability at the few % level and is potentially detectable. This nighttime probability starts with the same structure as for daytime solar neutrinos, except that instead of immediately projecting from the mass state to the flavor state at the detector, the mass state must propagate through the Earth to the detector. The nighttime probability is then
@@ IV.2 Galactic Supernova Neutrinos
The neutrino signal coming from a galactic supernova will contain a wealth of information about supernova dynamics and neutrino oscillations. Depending on the time of day, the signal may well travel through the Earth before detection which will modify the signal somewhat [94, 95, 96, 97] and the calculation of that probability will need a high degree of precision due to the importance of such a data set. This code can also be used for that by using the solar part but setting the production density much higher to quantify the MSW effect in the supernova and considering the fact that all six flavors are populated in the source.
@@ V Results
Figures 2-3 show the oscillograms for atmospheric neutrino oscillation probabilities as a function of energy and zenith angle in both disappearance and appearance channels, for neutrinos and antineutrinos, and for the normal and inverted mass orderings. These highlight the well known parametric resonance through the Earth’s core that depends on the mass ordering [99, 100]. The impact of non-zero detector depth can be seen in that for low energy neutrinos, $\lesssim$ few GeV, we see a deviation from $P_{\mu\mu}=1$ at $\cos\theta_{z}>0$.
@@
Figure 4: The $\nu_{e}$ disappearance probability from the center of the Sun to a detector in the Earth as a function of energy and zenith angle. The bands on the right show the annual average exposure for nighttime neutrinos at the latitudes for DUNE (blue) and SK/HK (red). The detector depth is at 2 km which leads to some effects for daytime neutrinos as well.
In fig. 5 we show the annual averaged nighttime probability for DUNE’s latitude along with the daytime probability with no terrestrial matter effect, clearly showing the effect of regeneration, see e.g. [102]. For further validation examples, see appendix A.
@@ VI Code Usage
$\bullet~\text{mixing angle}\,\theta_{23}~\bullet~\text{CP phase}\,\delta~\bullet\,\text{Production height}$
Any set of nested loops over a set of parameters that contains one or more of these three parameters should have these parameters on the innermost loops.
@@ VII.1 Speed
We benchmark the time it takes to compute one set of three eigenvalues and corresponding eigenvectors under the different compiler flags and for different levels of precision. We compute the time it takes to compute the eigenvalues and eigenvectors for a standard set of parameters 10,000 times to get an estimate of the time it takes to compute the eigenvalues once. We then repeat this 10,000 times to get an uncertainty due to variations in the computer’s state and also repeat the procedure for different compiler flags. Our results are shown in fig. 6 which shows that even at an eigenvalue precision of 2, which essentially reaches the double precision limit (see fig. 8 below), the approximation technique is still definitively faster than the exact expression from Cardano [22]. Our recommended precision is 1 (although 0 is also sufficient for all practical uses as well) as that is significantly more precise than will be needed for next generation atmospheric and solar measurements and provides a factor of $\sim 2$ speedup in the computing the eigenvalues and eigenvectors over the exact expression. We also recommend a compiler flag of -O3 as it is significantly faster than -O0 and -Ofast offers no real advantage with some additional compatibility issues.
@@
We combine everything into the total time required to compute one core crossing trajectory. We use the PREM_NUniformLayer with varying numbers of layers for simplicity. We compute the probabilities 10,000 times and take the mean time. We then repeat this 100 times to get an estimate of the mean and the uncertainty. The numerical results for different compiler flags are shown in fig. 7. The results are shown where either $\theta_{23}$ is changed for each of the 10,000 computations or $\Delta m^{2}_{31}$ is changed each time. We see that changing $\theta_{23}$ is faster at any Earth model and the cost to change it does not increase with the complexity of the Earth model, as it does for $\Delta m^{2}_{31}$. Similar results are achieved where changing $\delta$ or the production height also is extremely fast, while changing the other oscillation parameters, $\theta_{12}$, $\theta_{13}$, and $\Delta m^{2}_{21}$ require recomputing all the amplitudes through the Earth. This numerically shows that the so-called “fast” parameters, $\theta_{23}$, $\delta$, and the production height can be varied for essentially no significant computational cost.
@@
We now compare the speed of our code to other popular atmospheric neutrino oscillation codes in the literature. We benchmarked the speed of nuSQuIDS [47] across a range of energies and zenith angles. We noted that the code computed some trajectories ($\cos\theta_{z}$ closer to zero) much faster than others. We found that the average time to compute one probability was $1.6\times 10^{6}$ ns. We also note that while this calculation did not include new physics or neutrino interactions in the Earth, the nuSQuIDS code in general is designed to handle these scenarios. We also benchmarked OscProb [103] and found that for a core-crossing trajectory with 44 layers, the time to compute one probability is $\sim 40\,000$ ns, regardless of the oscillation parameter looped over, compared with $\sim 6000$ ns when looping over a slow parameter and $\sim 200$ ns when looping over a “fast” parameter.
@@ VII.2 Precision
We consider the precision in terms of the two main approximations used to compute atmospheric oscillation probabilities: the eigenvalues and the Earth density model. To estimate the error, we compute the oscillation probabilities across a $100\times 100$ grid of energies $E\in[2,40]$ GeV and zenith angles $\cos\theta_{z}\in[-1,0.1]$. We consider our “exact” oscillation probabilities as those with exact eigenvalues and one million total layers, spread out with 250,000 in each of the four main regions via the PREM_NDiscontinuityLayer (2.5e5, 2.5e5, 2.5e5, 2.5e5) functional call. We then recalculate the oscillation probabilities across this grid with approximate eigenvalues or fewer layers in the Earth model or with the two integration based Earth models: PREM_Full and PREM_Four. Then, to compare the two grids, we compute the maximum difference in the two main oscillation probabilities: $P(\nu_{\mu}\to\nu_{\mu})$ and $P(\nu_{\mu}\to\nu_{e})$ across this entire grid which covers the majority of the oscillation physics. Certainly using a finer (coarser) spaced grid would find more (fewer) points with a larger error, but the mean error remains relatively unchanged and the $100\times 100$ sized grid is reasonably fine grained for practical applications. Our results are summarized in fig. 8. We find that the mean error as defined above as a function of the total number of layers in the Earth model is well approximated by
@@
Figure 8: The mean error (blue square) and maximum error (orange arrow) in the probabilities across a $100\times 100$ grid in energy and zenith angles including both disappearance and appearance. The two approximations considered are shown: the eigenvalue precision, $N_{\rm NR}$, on the left and the mumber of layers in the Earth density model, $N_{\rm layer},$ on the right with different vertical scales. The integration methods PREM_Full and PREM_Four are also shown in the right panel.
We find that even with zero NR iterations, the eigenvalues are already incredibly precise introducing corrections no larger than $10^{-5}$ at the probability level, far below what can be measured. We also see that NR iterations to the eigenvalues yield vast further improvements in the probabilities all the way down to the double precision limit with just two iterations.
We see that the precision rapidly improves as the number of layers is increased. At 1000 total layers the maximum error is $10^{-3}$ and the mean error is $\sim 10^{-6}$ which is significantly more precise than will be needed in the coming experiments. In reality, $\sim 32$ total layers with sub-percent maximum error and $10^{-4}$ mean error is sufficient for nearly all practical uses. While this study kept the layers evenly divided among the four regions, further optimization there may be possible.
@@ VII.3 Recommendations
The two main choices to be made when using the code reflect the two approximations: the eigenvalues and the Earth density model.
@@
First we address the choices for the eigenvalues. While it may be tempting to simply set this to the exact expression (by calling any negative number in the Set_Eigenvalue_Precision function), this is not necessary. We see in fig. 8 that the difference between two NR and the exact expression is indistinguishable and fig. 6 shows that computing the eigenvalues and eigenvectors is over a third faster when using two NR iterations than when using the exact expression. Thus we never recommend using the exact Cardano expression. In fact, given the level of precision achieved in the approximations, we find it unlikely that an experiment can even differentiate between zero NR iterations and one NR iteration. Our official recommendation is to use one NR iteration and that is what the code uses by default, but zero NR iterations should not be distinguishable.
@@
Second is the more complicated question of the Earth models. Some have proposed using the integration technique implemented in the PREM_Full or the slightly simpler PREM_Four Earth density models. When comparing Earth models with a very small ($\lesssim 10$) layers, this approach may be beneficial. It addresses the problem of neutrinos just skimming into e.g. the outer core and seeing the average density of that region when the actual density seen is less due to the modest variation within that region. We find that the advantages of this approach do not outweigh the costs, which are significant. The initial integration time is slower (see table 2) but this is not likely to be significant. The main issues are twofold. First, we find that it introduces a correction compared to the model of using many steps. Second, it requires computing the eigenvalues and eigenvectors for each energy, each layer, and each zenith angle. So while the number of layers may be smaller, if the number of zenith angles is not small then it may be considerably more expensive then using many layers. Thus we recommend one of the constant density layered approaches. We have found that the uniform spacing of layers can lead to some problems near the sharp boundaries so we always recommend PREM_NDiscontinuityLayer over the similar (but also simpler) PREM_NUniformLayer. The number of layers needed does depend on the precision desired as shown in fig. 8. For nearly all applications PREM_NDiscontinuityLayer(2,10,10,5) should be sufficient. These numbers are chosen as the inner core does not vary too much and not many neutrinos pass through it (2), the outer core varies a lot (10), the inner mantle also has a sizable variation and experiences a large amount of the flux (10), and the outer mantle is a smaller region but with more fine grained features (5). If a high precision calculation is needed, increasing these layers by some factor (e.g. 10) should be suffice. Further optimization on the split among the regions may yield additional small improvements in speed and/or precision.
@@
Beyond the recommendations for the approximations, we again reiterate the most important recommendation, regardless of the choices made above. When looping over parameters, we strongly encourage that $\theta_{23}$, $\delta$, and the production height be in the innermost loops, if they are varied in a given calculation. This will lead to dramatic speed ups as nearly all of the calculations can be reused. The speed ups are the most dramatic for the most complicated Earth models as changing those parameters only require recomputing steps at the very beginning and/or the very end of the trajectory, but none of the interior steps.
@@ VIII Conclusions
Current generation neutrino oscillation experiments already require the computation of neutrino oscillation probabilities many times filling up super computers for days. Next generation experiments achieving higher statistics will require orders of magnitude more throws necessitating improved algorithms for computing neutrino oscillation probabilities efficiently. We have previously addressed this problem for long-baseline oscillations [18]. To handle the significantly more complicated problem of neutrinos from the atmosphere, the Sun, and supernova traversing the Earth with its sharply varying density profile requires a more sophisticated approach. Using some of the same neutrino theory and linear algebra results we efficiently propagate neutrinos through the Earth in a specific basis that does not depend on $\theta_{23}$ or $\delta$ applying these parameters only at the beginning and end of the trajectory. This and other features allow us to efficiently save and reuse computations as parameters are varied.
@@
  • The code intelligently saves often repeated calculations leading to significant speed ups. For example, nearly all of the calculation is unchanged when $\theta_{23}$ or $\delta$ changes; looping over those two variables is extremely fast. A similar speed up is also implemented for variations in the production height.

### v2

@@ IV.1 Solar Neutrinos
Nighttime solar neutrinos contain an important regeneration effect. It is known, based on oscillation probability parameters and some preliminary data, that the solar $\nu_{e}$ disappearance probability should be slightly higher, on average, at night than during the day. While the probability oscillates quite quickly, directly observing those oscillations requires energy resolution and statistics far beyond what is achievable. Nonetheless, the averaged effect modifies the probability at the few % level and is potentially detectable. This nighttime probability starts with the same structure as for daytime solar neutrinos, except that instead of immediately projecting from the mass state to the flavor state at the detector, the mass state must propagate through the Earth to the detector. The nighttime probability is then
@@ IV.2 Galactic Supernova Neutrinos
The neutrino signal coming from a galactic supernova will contain a wealth of information about supernova dynamics and neutrino oscillations. Depending on the time of day and location in the sky, the signal may well travel through the Earth before detection which will modify the signal somewhat [96, 97, 98, 99] and the calculation of that probability will need a high degree of precision due to the importance of such a dataset. This code can also be used for that by using the solar part but setting the production density much higher to quantify the MSW effect in the supernova. In addition, unlike for solar neutrinos which are essentially only electron neutrinos, all six flavors (including both neutrinos and antineutrinos) are populated in the source requiring the computation of all oscillation channels.
@@ V Results
Figures 2-3 show the oscillograms for atmospheric neutrino oscillation probabilities as a function of energy and zenith angle in both disappearance and appearance channels, for neutrinos and antineutrinos, and for the normal and inverted mass orderings. These highlight the well known parametric resonance through the Earth’s core that depends on the mass ordering [101, 102]. The impact of nonzero detector depth can be seen in that for low energy neutrinos, $\lesssim$ few GeV, we see a deviation from $P_{\mu\mu}=1$ at $\cos\theta_{z}>0$.
@@
We also investigated the role of detector depth by computing the difference in the oscillation probability for a detector on the surface and a detector 2 km below the surface. We summarized the results in table 2 which shows that, for atmospheric neutrinos, the variation in the probability is up to a few percent in some channels. Oscillograms of all 8 channels are included in the appendix for completeness, see Figs. 10-11. These show that, in most cases, the effect is for horizontal to slightly down-going trajectories at low energies, which is a fairly small fraction of phase space and does not contribute much to the oscillation information which tends to come from trajectories through the mantle and/or the core. Nonetheless, for precise calculations, including the nonzero detector depth is essential.
Table 2: Maximum shift in probability over $E\in[2,40]$ GeV and $\cos\theta_{z}\in[-1,0.1]$ between a detector on the surface and a detector at a depth of 2 km.
$\max|\Delta P|$ $\nu_{\mu}\to\nu_{\mu}$ $\nu_{\mu}\to\nu_{e}$
$\nu$ $\bar{\nu}$ $\nu$ $\bar{\nu}$
NO 5.1% 5.1% 1.5% 0.2%
IO 5.4% 5.4% 0.3% 1.5%
@@
Figure 4: The $\nu_{e}$ disappearance probability from the center of the Sun to a detector in the Earth as a function of energy and zenith angle. The bands on the right show the annual average exposure for nighttime neutrinos at the latitudes for DUNE (blue) and SK/HK (red). The detector depth is at 2 km which leads to some matter effects for daytime ($\cos\theta_{z}>0$) neutrinos as well.
In Fig. 5 we show the annual averaged nighttime probability for DUNE’s latitude along with the daytime probability with no terrestrial matter effect, clearly showing the effect of regeneration, see e.g. [104]. For further validation examples, see appendix B.
@@ VI Code Usage
$\bullet~\text{mixing angle}\,\theta_{23}~\bullet~\text{CP phase}\,\delta$
$\bullet\,\text{Production height}~\bullet\,\text{Solar density}$
Any set of nested loops over a set of parameters that contains one or more of these four parameters should have these parameters on the innermost loops.
@@ VII.1 Speed
We benchmark the time it takes to compute one set of three eigenvalues and corresponding eigenvectors under the different compiler flags and for different levels of precision. We compute the time it takes to compute the eigenvalues and eigenvectors for a standard set of parameters 10,000 times to get an estimate of the time it takes to compute the eigenvalues once. We then repeat this 10,000 times to get an uncertainty due to variations in the computer’s state. Finally, we repeat the procedure for different compiler flags. Our results are shown in Fig. 6 which shows that even at an eigenvalue precision of 2, which essentially reaches the double precision limit (see Fig. 9 below), the approximation technique is still definitively faster than the exact expression from Cardano [22]. Our recommended precision is 1 (although 0 is also sufficient for all practical uses as well) as that is significantly more precise than will be needed for next generation atmospheric and solar measurements and provides a factor of $\sim 2$ speedup in the computing the eigenvalues and eigenvectors over the exact expression. We also recommend a compiler flag of -O3 as it is significantly faster than -O0 and -Ofast offers no real advantage with some additional compatibility issues.
@@
We combine everything into the total time required to compute one core crossing trajectory. We use the PREM_NUniformLayer with varying numbers of layers for simplicity. We compute the probabilities 10,000 times and take the mean time. We then repeat this 100 times to get an estimate of the mean and the uncertainty. The numerical results for different compiler flags are shown in Fig. 7. The results are shown where either $\theta_{23}$ is changed for each of the 10,000 computations or $\Delta m^{2}_{31}$ is changed each time. We see that changing $\theta_{23}$ is faster at any Earth model and the cost to change it does not increase with the complexity of the Earth model, as it does for $\Delta m^{2}_{31}$. Similar fast results are achieved where changing $\delta$ or the production height, while changing the other oscillation parameters, $\theta_{12}$, $\theta_{13}$, and $\Delta m^{2}_{21}$ require recomputing all the amplitudes through the Earth. This numerically shows that the so-called “fast” parameters, $\theta_{23}$, $\delta$, and the production height can be varied for essentially no significant computational cost regardless of the complexity of the Earth model.
@@
The time to compute probabilities across many zenith angles and across many energies differs between these two parameters. Specifically, increasing the number of zenith angles grid points is cheaper than increasing the number of energy grid points with the difference being about a factor of 2, see Fig. 8. The difference is more modest for Earth models with $\lesssim 30$ layers before it asymptotes to this factor of $\sim 2$. We have confirmed that this trend holds quite generally regardless of the number of zenith angle or energy grid points, as well as the compiler options. The reason for this benefit is that the eigenvalues and eigenvectors need to be calculated for each layer in the Earth and for each density. Thus varying the zenith angle many times does not lead to an increase in the number of eigenvalue and eigenvector computations while varying the energy many times does. Both cases still require computing the amplitudes, multiplying them together, and then performing the final transformation from $\widetilde{\mathcal{A}}_{traj}$ to the probabilities, see Eqs. 36-37. What this result shows is that latter part of the computation takes a comparable amount of time as computing all the eigenvalues and eigenvectors.
Figure 8: The mean computation time for one core-crossing trajectory divided by the number of layers for different numbers of layers and for different compiler flags. The solid and dashed lines represent cases where the algorithm is used at 100 grid points in energies and zenith angles, respectively. Clearly, zenith angle scans are approximately twice as fast as energy scans using compiler flags -03 and -0fast.
@@
We now compare the speed of our code to other popular atmospheric neutrino oscillation codes in the literature. We benchmarked the speed of nuSQuIDS [48] across a range of energies and zenith angles. We noted that the code computed some trajectories ($\cos\theta_{z}$ closer to zero) much faster than others. We found that the average time to compute one probability was $1.6\times 10^{6}$ ns. We also note that while this calculation did not include new physics or neutrino interactions in the Earth, the nuSQuIDS code in general is designed to handle these scenarios. We also benchmarked OscProb [105] and found that for a core-crossing trajectory with 44 layers, the time to compute one probability is $\sim 40\,000$ ns, regardless of the oscillation parameter looped over, compared with $\sim 6000$ ns when looping over a slow parameter and $\sim 200$ ns when looping over a “fast” parameter in our NuFast-Earth code.
@@ VII.2 Precision
We consider the precision in terms of the two main approximations used to compute atmospheric oscillation probabilities: the eigenvalues and the Earth density model. To estimate the error, we compute the oscillation probabilities across a $100\times 100$ grid of energies $E\in[2,40]$ GeV and zenith angles $\cos\theta_{z}\in[-1,0.1]$. We consider our “exact” oscillation probabilities as those with exact eigenvalues and one million total layers, spread out with 250,000 in each of the four main regions via the PREM_NDiscontinuityLayer (2.5e5, 2.5e5, 2.5e5, 2.5e5) functional call. We then recalculate the oscillation probabilities across this grid with approximate eigenvalues or fewer layers in the Earth model or with the two integration based Earth models: PREM_Full and PREM_Four. Then, to compare the two grids, we compute the maximum and average difference across both of the two main oscillation probabilities: $P(\nu_{\mu}\to\nu_{\mu})$ and $P(\nu_{\mu}\to\nu_{e})$. The grid scanned is designed to cover the majority of the relevant oscillation physics for atmospheric neutrinos. Certainly using a finer (coarser) spaced grid would find more (fewer) points with a larger error, but the mean error remains relatively unchanged and the $100\times 100$ sized grid is reasonably fine grained for practical applications. Our results are summarized in Fig. 9. We find that the mean error as defined above as a function of the total number of layers in the Earth model is well approximated by
@@
Figure 9: The mean error (blue square) and maximum error (orange arrow) in the probabilities across a $100\times 100$ grid in energy and zenith angles including both disappearance and appearance. The two approximations considered are shown: the eigenvalue precision, $N_{\rm NR}$, on the left and the number of layers in the Earth density model, $N_{\rm layer},$ on the right with different vertical scales. The integration methods PREM_Full and PREM_Four are also shown in the right panel.
We see that the precision rapidly improves as the number of layers is increased. At 1000 total layers the maximum error is $10^{-3}$ and the mean error is $\sim 10^{-6}$ which is significantly more precise than will be needed in the coming experiments. In reality, $\sim 32$ total layers with subpercent maximum error and $10^{-4}$ mean error is sufficient for nearly all practical uses. While this study kept the layers evenly divided among the four regions, further optimization there may be possible.
@@
We find that even with zero NR iterations on the eigenvalues, the eigenvalues are already incredibly precise introducing corrections no larger than $10^{-5}$ at the probability level, far below what can be measured. We also see that NR iterations to the eigenvalues yield vast further improvements in the probabilities all the way down to the double precision limit with just two iterations.
@@ VII.3 Recommendations
There are a number of choices of how to implement neutrino oscillations through the Earth and we carefully go through our recommendations here.
@@
First we address the choice for the eigenvalues. While it may be tempting to simply set this to the exact expression (by calling any negative number in the Set_Eigenvalue_Precision function), this is not necessary. We see in Fig. 9 that the difference between two NR and the exact expression is indistinguishable and Fig. 6 shows that computing the eigenvalues and eigenvectors is over a third faster when using two NR iterations than when using the exact expression. Thus we never recommend using the exact Cardano expression. In fact, given the level of precision achieved in the approximations, we find it unlikely that an experiment can even differentiate between zero NR iterations and one NR iteration. Our official recommendation is to use one NR iteration and that is what the code uses by default, but zero NR iterations should not be distinguishable and is slightly faster.
@@
Second is the more complicated question of the Earth models. Some have proposed using the integration technique implemented in the PREM_Full or the slightly simpler PREM_Four Earth density models. When comparing Earth models with a very small ($\lesssim 10$) layers, this approach may be beneficial. It addresses the problem of neutrinos just skimming into e.g. the outer core and seeing the average density of that region when the actual density seen is less due to the modest variation within that region. We find that the advantages of this approach do not outweigh the costs, which are significant. The initial integration time is slower (see table 3) but this is not likely to be significant. The main issues are twofold. First, we find that it introduces a correction compared to the model of using many steps. Second, it requires computing the eigenvalues and eigenvectors for each energy, each layer, and each zenith angle. So while the number of layers may be smaller, if the number of zenith angles is not small then it may be considerably more expensive than using many layers. Thus we recommend one of the constant density layered approaches. We have found that the uniform spacing of layers can lead to some problems near the sharp boundaries so we always recommend PREM_NDiscontinuityLayer over the similar (but also simpler) PREM_NUniformLayer. The number of layers needed does depend on the precision desired as shown in Fig. 9. For many applications PREM_NDiscontinuityLayer(2,10,10,5) should be sufficient. These numbers are chosen as the inner core does not vary too much and not many neutrinos pass through it (2), the outer core varies a lot (10), the inner mantle also has a sizable variation and experiences a large amount of the flux (10), and the outer mantle is a smaller region but with more fine grained features (5). If a high precision calculation is needed, increasing these layers by some factor (e.g. 10) should be suffice. Further optimization on the split among the regions may yield additional small improvements in speed and/or precision.
@@
Beyond the recommendations for the approximations, we again reiterate the most important recommendation, regardless of the choices made above. When looping over parameters, we strongly encourage that $\theta_{23}$, $\delta$, the production height, and the density in the Sun, be in the innermost loops, if they are varied in a given calculation. This will lead to dramatic speed ups as nearly all of the calculations can be reused. The speed ups are the most dramatic for the most complicated Earth models as changing those parameters only require recomputing steps at the very beginning and/or the very end of the trajectory, but none of the interior steps and thus do not depend on the number of Earth layers. Even if this is not possible, e.g. in a Markov chain approach, some steps in the fast directions will still be fast, as will the integrals over the production height (atmospheric) or production region (solar).
We note that the atmosphere is extremely well approximated as vacuum, an approximation which provides considerable speed up as any integral over production height becomes essentially free. Locating the detector at the correct detector depth is important as this can shift the probability $1\%$, especially for horizontal trajectories.
Finally, we comment on energy versus zenith angles. Since the eigenvalues and eigenvectors must be recomputed for each layer and each energy, a large number of energies requires many computations of the eigenvalues and eigenvectors. If a comparable overall precision can be achieved via interpolation with a shift to more zenith angles and fewer energies in the array, then this can save computational time as described in Fig. 8.
@@ VIII Conclusions
Current generation neutrino oscillation experiments already require the computation of neutrino oscillation probabilities many times filling up super computers for days. Next generation experiments achieving higher statistics will require orders of magnitude more throws necessitating improved algorithms for computing neutrino oscillation probabilities efficiently. We have previously addressed this problem for long-baseline oscillations [18]. To handle the significantly more complicated problem of neutrinos traversing the Earth with its sharply varying density profile coming from the atmosphere, the Sun, and a supernova, requires a more sophisticated approach. Using some of the same neutrino theory and linear algebra results along with new components necessary to address the Earth’s sharply varying density profile, we efficiently propagate neutrinos through the Earth in a specific basis that does not depend on $\theta_{23}$ or $\delta$ applying these parameters only at the beginning and end of the trajectory. This and other features allow us to efficiently save and reuse computations as parameters are varied.
@@
  • The code intelligently saves often repeated calculations leading to significant speed ups. For example, nearly all of the calculation is unchanged when $\theta_{23}$ or $\delta$ changes; looping over those two variables is extremely fast. A similar speed up is also implemented for variations in the production height (atmospheric) or density in the Sun (solar neutrinos).
  • Eigenvalue and eigenvector calculations are also cached for each energy and Earth layer, allowing for rapid computations over zenith angle.